A 7500-kg rocket blasts off vertically from the launch pad with a constant upward acceleration of 2.25 m/s2 and feels no appreciable air resistance. When it has reached a height of 525 m, its engines suddenly fail so that the only force acting on it is now gravity. (a) What is the maximum height this rocket will reach above the launch pad? (b) How much time after engine failure will elapse before the rocket comes crashing down to the launch pad, and how fast will it be moving just before it crashes? (c) Sketch ay-t, vy-t, and y-t graphs of the rocket’s motion from the instant of blast-off to the instant just before it strikes the launch pad.


1
Expert's answer
2021-02-28T07:27:36-0500

a)


"v_y=\sqrt{2ah}=\sqrt{2(2.25)(525)}=48.6\frac{m}{s}"

"y=h+\frac{v_y^2}{2g}=525+\frac{48.6^2}{2(9.8)}=645.5\ m"

b)


"h=-v_yt+0.5gt^2\\525=-48.6t+4.9t^2\\t=16.44\ s"

"V_y^2=v_y^2+2gh\\V_y^2=48.6^2+2(9.8)(525)\\V_y=112.5\frac{m}{s}"

c)